OpenStudy (anonymous):

What is i^-9 ( That's i raised to the negative ninth) Not the variable i, i as in an imaginary number.

OpenStudy (anonymous):

\[i^{-2}=\frac{ 1 }{ -1 }\] \[i^{-1}=\frac{1}{\sqrt{-1}}\] \[i^{-3}=\frac{1}{(-1)^{3/2}}\]

OpenStudy (anonymous):

combine these identities until you reach i^(-9)